Problom 2.5.4. (Price of continuous random dividends) The current price of a stock is 100. The stock pays dividends continuously at a rate proportional to its price. The dividend yields is 396. The continuously compounded risk-free interest rate is 7%. Calculate the price of the stream of dividends to be paid in the next 5 years. (Note: Because the dividends are stochastic, their present value is also stochastic and hence cannot be their price.)
Answer #1

Answers

Say that an investor longs one share of stock. Then, with continuous reinvestment of dividends the investor’s profit can be expressed, in our usual notation, as S(T)e δT − S(0)e rT .

A rational investor who demands to be compensated for risk would only invest if the expected profit above were positive.

So, X = E[S(T)] > S(0)e (r−δ)T = 100e (0.07-0.03)5=100e(0.04)5=100e(0.2)= 100(1.2214)=122.14(Answer)

r= risk free rate=7%1

δ=dividend yield =3%

S= spot price

T= time period = 5 years

x= price